At the local convenience store, 6 bags of chips and 4 containers of dip cost $34. However, 4 bags of chips and 2 containers of dip cost $20. What is the cost of one bag of chips and one container of dip?
Answer by mathstutor494(120)

Let "c" and "d" be cost per unit bag of chips and container of dip respectively
Given in the problem.... 6 bags of chips and 4 containers of dip cost $34
which is expressed as
6c+4d=34............(1)
Also given that 4 bags of chips and 2 containers of dip cost $20
which is expressed as
4c+2d=20........(2)
Multiplying eqn no. 2 by 2 results in
8c+4d=40..........(3)
Subtracting eqn no 1 from 3 results in
(8c+4d)-(6c+4d)=40-34
8c+4d-6c-4d=6
2c=6 ( as 4d gets cancelled out)
dividing both sides by 2
c=3
Now substituting value of c in eqn no.1 to get value of d.....
6*3+4d=34
18+4d=34
Subtracting 18 from either side in above equation results in
4d=16
Dividing both sides by 4
d=4
Ans : cost of one bag of chips $3 and one container of dip $4